4479:

A goldsmith melts 12.4 grams of gold to make a ring. The temperature of the gold rises from 26°C to 1064°C, and then the gold melts completely. If gold’s specific heat is 0.1291 joules/gram degree Celsius and its heat of fusion is 63.5 joules/gram, how much energy is gained by the gold? The gold gains a total of _________joules of energy.

PoseidonGodOfTheSea:

Problem One You will use both m * c * deltaT and H = m * heat of fusion. Givens m = 12.4 grams c = 0.1291 t1 = 26oC t2 = 1204 heat of fusion (H_f) = 63.5 J/grams. Equation H = m * c * deltaT + m * H_f Solution H = 12.4 * 0.1291 * (1063 - 26) + 12.4 * 63.5 H = 1660.1 + 787.4 H = 2447.5 or 2447.47 is the exact answer. I have to leave the rounding to you. I have no idea where to round it although I suspect 2450 would be right for 3 sig digs.

PoseidonGodOfTheSea:

Problem Two Formula and Givens t1 = 14.5 t2 = 50.0 E = 5680 c = 4.186 m = ?? E = m c * deltaT Solution 5680 = m * 4.186 * (50 - 14.5) 5680 = m * 4.186 * (35.5) 5680 = m * 148.603 * m m = 5680 / 148.603 m = 38.22 grams That isn't very much. Be very sure you are working in joules. You'd leave that many grams in the kettle after drying it thoroughly. m = 38.2 to 3 sig digs.

mymissingsoul:

Copy and pasted wow

4479:

O_O

PoseidonGodOfTheSea:

I have done this question before I use documents

PoseidonGodOfTheSea:

and so yes I copied and pasted my own work now leave me alone

4479:

The gold gains a total of _________joules of energy.

mymissingsoul:

Lies I found your source on brainly

mymissingsoul:

it's 2448 by the way.

PoseidonGodOfTheSea:

brainly? hell I dont use brainly

PoseidonGodOfTheSea:

moron

4479:

o

@mymissingsoul wrote:
it's 2448 by the way.

4479:

@poseidongodofthesea wrote:
moron
report.

mymissingsoul:

@poseidongodofthesea wrote:
brainly? hell I dont use brainly
Liar want me to expose your source for an answer

mymissingsoul:

Problem One You will use both m * c * deltaT and H = m * heat of fusion. Givens m = 12.4 grams c = 0.1291 t1 = 26oC t2 = 1204 heat of fusion (H_f) = 63.5 J/grams. Equation H = m * c * deltaT + m * H_f Solution H = 12.4 * 0.1291 * (1063 - 26) + 12.4 * 63.5 H = 1660.1 + 787.4 H = 2447.5 or 2447.47 is the exact answer. I have to leave the rounding to you. I have no idea where to round it although I suspect 2450 would be right for 3 sig digs. Problem Two Formula and Givens t1 = 14.5 t2 = 50.0 E = 5680 c = 4.186 m = ?? E = m c * deltaT Solution 5680 = m * 4.186 * (50 - 14.5) 5680 = m * 4.186 * (35.5) 5680 = m * 148.603 * m m = 5680 / 148.603 m = 38.22 grams That isn't very much. Be very sure you are working in joules. You'd leave that many grams in the kettle after drying it thoroughly. m = 38.2 to 3 sig digs.

PoseidonGodOfTheSea:

yes I do cause I know that I dont use brainly I use https://answer-ya.com/ as a homework help site as well

mymissingsoul:

His source.

4479:

=_=

mymissingsoul:

Smh

PoseidonGodOfTheSea:

lel bro you didnt do anything all you did was say the exact same thing as me

PoseidonGodOfTheSea:

and that is my work ik the answer

mymissingsoul:

I didn't. I just gave the numbers 2448.

PoseidonGodOfTheSea:

because? I had help on the same question from a student on the site

mymissingsoul:

And you didn't work anything out liar.

PoseidonGodOfTheSea:

yes I did now LEAVE ME ALONE YOU CYBER BULLY THATS ALL YOU DO LITERALLY

mymissingsoul:

Lol imagine not doing your work and lying about doing it

PoseidonGodOfTheSea:

I did my work and I never lied

mymissingsoul:

copy and paste isn't your work